# A consistent, volume preserving, and adaptive mesh refinement-based framework for modeling non-isothermal gas-liquid-solid flows with phase change

### v1

@@ 7 Results and discussion
7.1 Isothermal advection of a dense bubble
To demonstrate the combined consistency of mass, momentum, and enthalpy integrators, we modify the dense bubble advection problem discussed in Sec. 3.2: in addition to mass and momentum transport, we also consider enthalpy transported by the bubble. Thermal conductivity $\kappa$ and $Q_{\text{src}}$ are set to zero. All other problem parameters (e.g., bubble size, boundary conditions, etc.) remain the same as in Sec. 3.2. In the absence of thermal diffusion and body forces, the system of equations reduces to
$\frac{\partial\phi}{\partial t}+\bm{\mathrm{u}}\cdot\nabla{\phi}=0,$ (55a)
@@ 7.1 Isothermal advection of a dense bubble
    The bubble gets advected with constant velocity, retains its shape, and returns to its original position after each time period.
@@
Fig. 8 illustrates the advection of the dense bubble at various time instances. The simulation employs $l=2$ levels of mesh with the coarse grid size set to $N_{x0}\times N_{y0}=128^{2}$, and a refinement ratio of $n_{\text{ref}}=2$. The simulation runs until $t=10$ s. As shown in Fig. 8, the bubble moves with a constant velocity and returns to its original position after one time period $\tau=\frac{d}{|\bm{\mathrm{u}}_{c}|}=1$ , in which $d=\sqrt{2}$ and $|\bm{\mathrm{u}}_{c}|=\sqrt{2}$. Additionally, the bubble undergoes minimal distortion while returning to its original position. Throughout the simulation (data not shown here for brevity), velocity remains constant at its initial value $\bm{\mathrm{u}}_{c}=(u,v)=(1,1)$. In addition, the enthalpy equation solution does not result in spurious phases (for example, liquid becoming solid).
@@
To quantify the accuracy of consistent mass-momentum-enthalpy integrators, a grid convergence study is conducted. We consider uniform grids of size $N_{x}\times N_{y}=\{128^{2},256^{2},512^{2}\}$ for the convergence study. This is to avoid discretization errors near the coarse-fine boundary of adaptively refined grids. Errors in total mass $\mathcal{M}=\int_{\Omega}\rho\,\rm{d}V$, $x$-momentum $\mathcal{L}_{x}=\int_{\Omega}\rho u\,\rm{d}V$, $y$-momentum $\mathcal{L}_{y}=\int_{\Omega}\rho v\,\rm{d}V$ and enthalpy $\mathcal{H}=\int_{\Omega}\rho h\,\rm{d}V$ are computed using Eq. (42), and the results are shown in Fig. 9. The system’s mass, momentum, and enthalpy remain constant over time, with the exception of the initial transient phase where the initially sharp interface diffuses. Additionally, the temperature remains mostly constant throughout the simulation ($T\approx 2$ K) except near the interface (data not shown for brevity). This is because the primary variable $h$ diffuses over time and $T$ is reconstructed based on $T$-$h$ relations in our solution algorithm. The temperature change near the interface does not induce phase change, however. Mass and momentum are at least third-order accurate ($\mathcal{O}(\Delta x^{3})$) using this method. As velocity remains constant (numerically), mass and momentum errors are the same. However, enthalpy solution is only first-order ($\mathcal{O}(\Delta x)$) accurate. The behavior of error as a function of grid resolution can be explained as follows. At the end of each time step, the density field is synchronized with the level set function. This limits density (and momentum, in this case) diffusion due to advection. The level set function is reinitialized to a signed distance function using a third-order accurate ENO scheme; see our prior works on level set reinitialization [29]. Enthalpy transport, however, is subject to the upwinding errors of the CUI limiter, which displays first-order accuracy near non-monotone regions. These are regions close to the liquid-gas interface. The CUI scheme’s first-order accuracy corresponds to Gudonov’s order barrier theorem.
@@
Figure 8: Advection of a dense liquid bubble with density ratio $\rho_{i}/\rho_{o}=10,000$ in an initially uniform velocity field $\bm{\mathrm{u}}=(u,v)=(1,1)$. The simulation is performed using two levels of mesh refinement. The coarse grid size is $N_{x}\times N_{y}=128^{2}$. A unit periodic domain is considered. $\kappa$, $\mu$ and $Q_{\text{src}}$ are set to zero.
@@ 7.2 Thermocapillary flows
in which $\sigma_{0}$ and $\frac{\rm{d}\sigma}{\rm{d}T}\big{|}_{T_{0}}$ are the surface tension and Marangoni coefficients (respectively) computed at the reference temperature $T_{0}$. The gas bubble moves upward due to the variation in surface tension caused by the imposed temperature gradient. Key non-dimensional parameters of the problem are the Reynolds number ${Re=\frac{\rho^{\rm L}VR}{\mu^{\text{L}}}}$, Marangoni number ${Ma=\frac{VR}{\alpha^{\text{L}}}}$, Capillary number ${Ca=\frac{V\mu^{\text{L}}}{\sigma_{0}}}$, and Prandtl number ${Pr=\frac{\mu^{\text{L}}C^{\rm L}}{\kappa^{\rm L}}}$. Here, $V$ is the velocity scale defined as ${V=\bigg{|}\frac{\frac{\rm{d}\sigma}{\rm{d}T}\big{|}_{T_{0}}\nabla TR}{\mu^{\text{L}}}}\bigg{|}$. Two different cases are considered: zero Marangoni and finite Marangoni. The zero Marangoni case checks the accuracy of the surface tension force model. The finite Marangoni case examines the low Mach enthalpy integrator’s ability to simulate thermocapillary flows.
@@ 7.2.2 Finite Marangoni number
The simulations are run until $t=0.12$ s. Grid convergence studies are performed in both 2D and 3D. The 2D grids considered have coarse level resolutions of $N_{0}=64,128$, and $256$ with two levels of mesh refinement and $n_{\text{ref}}=2$. A uniform time step size of $\Delta t=2.5\times 10^{-5}$ s is used for the coarse mesh ($N_{0}=64$). This choice of $\Delta t$ respects the explicit capillary time step size restriction [27, 47], which is given by $\Delta t_{\sigma}=\sqrt{\frac{\rho^{\rm L}+\rho^{\rm G}}{4\pi\sigma}\Delta x^{3}}\approx 2.33\times 10^{-5}$ s. With each successive grid refinement (by a factor of 2), the time step size is reduced by a factor of 4  [^2] Compared to the theoretical formula, this reduction seems more conservative, which would reduce $\Delta t$ by a factor of about $\sqrt{2^{3}}\approx 3$. In practice, a large variation in the curvature of the interface further reduces $\Delta t$.. For 3D simulations, coarse level grids of size $N_{0}=32,64$, and $128$ are considered, with a uniform time step size of $\Delta t=2\times 10^{-5}$ s for the coarse mesh. $\Delta t$ is halved for each fine grid, which maintains the explicit capillary time step size restriction.
@@ 7.3.1 Two-phase melting problem with flux boundary condition
Consider a quasi one-dimensional computational domain $\Omega\in[0,0.1]\times[0,0.005]$ as illustrated in Fig. 13(A). The simulation employs a two-level mesh with coarse cell size of $\Delta x=\Delta y=0.005/32$. The simulation runs until $t=10$ s with a constant time step size of $\Delta t=10^{-3}$ s. Initially, the domain is fully occupied by the solid PCM with a temperature of $T_{i}=1500$ K. At the right boundary ($x=0.1$), a zero pressure (outflow) condition is used for the flow solver and an adiabatic (homogeneous Neumann) condition is used for the temperature $T$. At the left end of the domain ($x=0$), no-slip and inhomogeneous Neumann ${\left.-\kappa^{\rm L}\frac{{\rm d}T^{\rm L}_{\text{exact}}}{{\rm d}x}\right|_{x=0}}$ boundary conditions are employed for velocity and temperature, respectively. Here, $T^{\rm L}_{\text{exact}}$ denotes the analytical temperature of the liquid phase as derived in A. The top and bottom boundaries (in the $y$-direction) are periodic.
The schematic of the problem is depicted in Fig. 13(A). Since $R_{\rho}=\rho^{\rm S}/\rho^{\rm L}>1$, there is a volume expansion upon melting, which induces additional flow. As shown in A, the solid phase moves with a rigid body velocity $u^{\rm S}(t)$, while the liquid velocity $u^{\rm L}$ remains zero. To simulate this behavior, the solid fraction $\varphi_{\rm S}$ in the Carman-Kozeny drag coefficient is replaced with the liquid fraction $\varphi$. Consequently, the Carman-Kozeny drag force retards liquid phase motion. Gradient-based tagging is used to generate adaptively refined grids. Melting initiates due to the flux boundary condition and the liquid-solid interface moves rightwards over time until the PCM completely melts. The evolution of the liquid fraction and the adaptive mesh is shown in the left column of Fig. 13. We compare the numerical solutions (interface position $x*=s(t)$, velocity of the solid $u^{\rm S}(t)$, temperature distribution $T(x)$) with the analytical ones in Fig. 14. An excellent match is observed between the two. The new analytical and standard Stefan problem solutions are very similar for this problem. This is attributed to the 1D nature of the problem as well as the low density difference between the solid and liquid phases ($\approx 11\%$). Nevertheless, this amount of density variation may cause PCM free surfaces to rise or protrude in 2D or 3D problems [1].
@@
Figure 13: Stefan problem with melting: time evolution of the liquid fraction $\varphi$ and adaptive grids. For the left column plots, melting occurs due to the flux boundary condition at the left end of the domain. For the right column plots, flux boundary conditions are imposed through a volumetric term in the enthalpy equation. There is also a passive gas phase. Gradient-based tagging is used to generate AMR grids. Solid, liquid and gas regions are depicted in yellow, blue and white colors.
@@ 7.3.2 Three-phase melting problem with a volumetric heat source
We now consider the three-phase version of the melting problem in which the heat flux boundary condition is imposed through a volumetric source term $Q_{\text{src}}$. The position of the heat source remains fixed at $x=0$, the initial location of the gas-PCM interface. Over time, the heat source strength varies (reduces). The PCM thermophysical properties remain the same as in the two-phase case.
@@
Due to the heat source, melting begins and progresses towards the right, as shown in Fig. 13. Comparison between numerical and analytical solutions is presented in Fig. 15. The numerical and analytical solutions are in excellent agreement. Thus, this simple test problem facilitates the validation of three phase codes aimed at modeling laser-induced PCM melting.
@@ 7.4 Porosity defect modeling
The final test case involves liquid, solid, and gas interactions. The goal is to simulate porosity defects during metal solidification, such as during metal 3D printing and casting. We consider a square computational domain $\Omega\in[0,L]^{2}$ of length $L=10^{-2}$ m. Fig. 16 depicts the initial phase distribution. The solid phase of the PCM occupies the region up to $y=0.1L$, the liquid phase of the PCM extends above it to $y=0.75L$, and a gas phase occupies the remainder. Three gas bubbles are placed within the liquid phase. The locations and radii of these bubbles can be found in Table 3.
@@
The simulation employs a two-level mesh with $N_{x0}\times N_{y0}=128^{2}$ coarse level cells, and a refinement ratio of 2 ($n_{\text{ref}}=2$). It runs until $t=1$ s with a uniform time step size of $3\times 10^{-6}$ s, which is determined by the explicit capillary time step size constraint. The PCM-gas interfacial region is refined using value-based ($|\phi|_{i,j}\leq 2\Delta x_{0}$) tagging criterion. Gradient-based tagging refines the solid-liquid interfacial regions. As bubbles rise due to buoyancy, solidification occurs concurrently at the bottom wall due to the imposed temperature boundary condition. The movement of bubbles induces convective flow in the liquid below, which helps solidify it more quickly. After a while, the liquid-solid interface entraps the left and right bubbles, while the center bubble breaks off from the liquid-gas interface and enters the gas phase. Due to surface tension, the breaking of the liquid-gas interface initiates capillary waves, which eventually dissipate due to the combined effects of fluid viscosity and freezing. In the meantime, the two remaining bubbles are trapped within the solid. As a result of the slow solidification rate above the bubbles and the motion of the remaining untrapped portion, the left and right bubbles take on the shape of a pancake (flat on the top). Solidification continues until the liquid PCM fully solidifies. This is a shrinkage problem ($R_{\rho}=\rho^{\rm S}/\rho^{\rm L}>1$), so the top PCM-gas interface moves downwards due to volume changes, as shown in final Fig. 16(E). Upon completion of solidification, the interfaces remain stationary. At this point, the liquid phase has completely disappeared from the domain. The results of the previous Stefan problem section, as well as this section highlight the strength of our method to (i) handle the appearance and disappearance of phases in the simulation; (ii) properly account for volume changes in the material; and (iii) maintain equilibrium with no spurious phase generation in the absence of a phase change triggering mechanism.
@@
Furthermore, we quantify the mass/volume conservation of our scheme in this case. We rerun the simulation after removing the center bubble from the domain to obtain meaningful convergence rates  [^3] Convergence rates are well-defined for smooth functions/dynamics.. This prevents interface breaking and keeps dynamics smooth. Fig. 17 shows the evolution of the interfaces. The solidification and bubble trapping dynamics are very similar to the three-bubble case discussed above. Uniform grids are used to study this problem. The PCM mass is computed numerically as
@@
The relative change in the PCM mass is obtained using Eq. (42) and plotted against time for four uniform grids of size $N_{x}\times N_{y}=\{64^{2},128^{2},256^{2},512^{2}\}$ in Fig. 18. The coarsest grid $(N_{x}\times N_{y}=64^{2})$ uses a uniform time step size of $\Delta t=3\times 10^{-5}$ s. For each successively refined grid, the time step size is reduced by a factor of 3, respecting the explicit capillary time step size restriction. It can be observed that as the grid cell size $\Delta$ decreases, the error in mass change of PCM, $\mathcal{E_{M}}$, also decreases till $t=0.3$ s. Around $t=0.3$ s, the interface between gas-PCM and liquid-solid comes close to one another, resulting in an error in mass calculation because two types of diffuse interfaces (solid-liquid interface and gas-PCM interface) meet and equilibrate rapidly, causing the PCM to solidify completely. However, the % mass change of the PCM for all grids is quite low ($\approx 0.2\%$). The relative change in bubbles’ volume is illustrated in Fig. 19. For all grids considered, the maximum volume change for the two bubbles is 2%. Given that the level set method leads to mass/volume loss particularly at coarse grid resolutions, this is an acceptable level of volume change.

### v2

@@ 7 Results and discussion
7.1 Isothermal advection of a droplet
To demonstrate the combined consistency of mass, momentum, and enthalpy integrators, we modify the droplet advection problem discussed in Sec. 3.2: in addition to mass and momentum transport, we also consider enthalpy transported by the droplet [^3] This benchmark test is provided in IBAMR GitHub within the directory examples/phase_change/ex2.. Thermal conductivity $\kappa$ and $Q_{\text{src}}$ are set to zero. All other problem parameters (e.g., droplet size, boundary conditions, etc.) remain the same as in Sec. 3.2. In the absence of thermal diffusion and body forces, the system of equations reduces to
$\frac{\partial{\color[rgb]{0,0,0}\Phi}}{\partial t}+\bm{\mathrm{u}}\cdot\nabla{{\color[rgb]{0,0,0}\Phi}}=0,$ (56a)
@@ 7.1 Isothermal advection of a droplet
    The droplet gets advected with constant velocity, retains its shape, and returns to its original position after each time period.
@@
Fig. 9 illustrates the advection of the droplet at various time instances. The simulation employs $l=2$ levels of mesh with the coarse grid size set to $N_{x0}\times N_{y0}=128^{2}$, and a refinement ratio of $n_{\text{ref}}=2$. The simulation runs until $t=10$ s. As shown in Fig. 9, the droplet moves with a constant velocity and returns to its original position after one time period $\tau=\frac{d}{|\bm{\mathrm{u}}_{c}|}=1$ , in which $d=\sqrt{2}$ and $|\bm{\mathrm{u}}_{c}|=\sqrt{2}$. Additionally, the droplet undergoes minimal distortion while returning to its original position. Throughout the simulation (data not shown here for brevity), velocity remains constant at its initial value $\bm{\mathrm{u}}_{c}=(u,v)=(1,1)$. In addition, the enthalpy equation solution does not result in spurious phases (for example, liquid becoming solid).
@@
To quantify the accuracy of consistent mass-momentum-enthalpy integrators, a grid convergence study is conducted. We consider uniform grids of size $N_{x}\times N_{y}=\{128^{2},256^{2},512^{2}\}$ for the convergence study. This is to avoid discretization errors near the coarse-fine boundary of adaptively refined grids. Errors in total mass $\mathcal{M}=\int_{\Omega}\rho\,\rm{d}V$, $x$-momentum $\mathcal{L}_{x}=\int_{\Omega}\rho u\,\rm{d}V$, $y$-momentum $\mathcal{L}_{y}=\int_{\Omega}\rho v\,\rm{d}V$ and enthalpy $\mathcal{H}=\int_{\Omega}\rho h\,\rm{d}V$ are computed using Eq. (43), and the results are shown in Fig. 10. The system’s mass, momentum, and enthalpy remain constant over time, with the exception of the initial transient phase where the initially sharp interface diffuses. Additionally, the temperature remains mostly constant throughout the simulation ($T\approx 2$ K) except near the interface (data not shown for brevity). This is because the primary variable $h$ diffuses over time and $T$ is reconstructed based on $T$-$h$ relations in our solution algorithm. The temperature change near the interface does not induce phase change, however. As shown in the Fig. 11, the mass and momentum transport is third-order accurate ($\mathcal{O}(\Delta x^{3})$) using this method. As velocity remains constant (numerically), mass and momentum errors are the same. However, enthalpy solution is only first-order ($\mathcal{O}(\Delta x)$) accurate. The behavior of error as a function of grid resolution can be explained as follows. At the end of each time step, the density field is synchronized with the level set function. This limits density (and momentum, in this case) diffusion due to advection. The level set function is reinitialized to a signed distance function using a third-order accurate ENO scheme; see our prior works on level set reinitialization [29]. Enthalpy transport, however, is subject to the upwinding errors of the CUI limiter, which displays first-order accuracy near non-monotone regions. These are regions close to the liquid-gas interface. The CUI scheme’s first-order accuracy corresponds to Gudonov’s order barrier theorem.
@@
Figure 9: Advection of a droplet with density ratio $\rho_{i}/\rho_{o}=10,000$ in an initially uniform velocity field $\bm{\mathrm{u}}=(u,v)=(1,1)$. The simulation is performed using two levels of mesh refinement. The coarse grid size is $N_{x}\times N_{y}=128^{2}$. A unit periodic domain is considered. $\kappa$, $\mu$ and $Q_{\text{src}}$ are set to zero.
@@
Figure 11: Convergence rates of the low Mach enthalpy method considering the isothermal advection of a droplet. The $\mathcal{L}^{2}$ error for a quantity $\psi$ is defined to be the root mean squared error (RMSE) of the vector $||\mathcal{E}_{\psi}||_{\rm RMSE}=||\mathcal{E}_{\psi}||_{2}/\sqrt{\mathcal{N}}$, in which $\mathcal{N}$ denotes the size of the vector $\mathcal{E}_{\psi}$. Here, $\Psi$ represents mass $\mathcal{M}=\int_{\Omega}\rho\,\rm{d}V$, $x$-momentum $\mathcal{L}_{\bm{\mathrm{x}}}=\int_{\Omega}\rho u\,\rm{d}V$, $y$-momentum $\mathcal{L}_{\bm{\mathrm{x}}}=\int_{\Omega}\rho v\,\rm{d}V$ and enthalpy $\mathcal{H}=\int_{\Omega}\rho h\,\rm{d}V$.
@@ 7.2 Thermocapillary flows
in which $\sigma_{0}$ and $\frac{\rm{d}\sigma}{\rm{d}T}\big{|}_{T_{0}}$ are the surface tension and Marangoni coefficients (respectively) computed at the reference temperature $T_{0}$. The gas bubble moves upward due to the variation in surface tension caused by the imposed temperature gradient. Key non-dimensional parameters of the problem are the Reynolds number ${Re=\frac{\rho^{\rm L}VR}{\mu^{\text{L}}}}$, Marangoni number ${Ma=\frac{VR}{\alpha^{\text{L}}}}$, Capillary number ${Ca=\frac{V\mu^{\text{L}}}{\sigma_{0}}}$, and Prandtl number ${Pr=\frac{\mu^{\text{L}}C^{\rm L}}{\kappa^{\rm L}}}$. Here, $V$ is the velocity scale defined as ${V=\bigg{|}\frac{\frac{\rm{d}\sigma}{\rm{d}T}\big{|}_{T_{0}}\nabla TR}{\mu^{\text{L}}}}\bigg{|}$. Two different cases are considered: zero Marangoni and finite Marangoni. The zero Marangoni case checks the accuracy of the surface tension force model [^5] This benchmark test is provided in IBAMR GitHub within the directory examples/multiphase_flow/ex14.. The finite Marangoni case examines the low Mach enthalpy integrator’s ability to simulate thermocapillary flows [^6] This benchmark test is provided in IBAMR GitHub within the directory examples/phase_change/ex7..
@@ 7.2.2 Finite Marangoni number
The simulations are run until $t=0.12$ s. Grid convergence studies are performed in both 2D and 3D. The 2D grids considered have coarse level resolutions of $N_{0}=64,128$, and $256$ with two levels of mesh refinement and $n_{\text{ref}}=2$. A uniform time step size of $\Delta t=2.5\times 10^{-5}$ s is used for the coarse mesh ($N_{0}=64$). This choice of $\Delta t$ respects the explicit capillary time step size restriction [31, 52], which is given by $\Delta t_{\sigma}=\sqrt{\frac{\rho^{\rm L}+\rho^{\rm G}}{4\pi\sigma}\Delta x^{3}}\approx 2.33\times 10^{-5}$ s. With each successive grid refinement (by a factor of 2), the time step size is reduced by a factor of 4  [^7] Compared to the theoretical formula, which would reduce $\Delta t$ by a factor of about $\sqrt{2^{3}}\approx 3$, a reduction of 4 seems more conservative. In practice, a large variation in the curvature of the interface further reduces $\Delta t$.. For 3D simulations, coarse level grids of size $N_{0}=32,64$, and $128$ are considered, with a uniform time step size of $\Delta t=2\times 10^{-5}$ s for the coarse mesh. $\Delta t$ is halved for each fine grid, which maintains the explicit capillary time step size restriction.
@@ 7.3.1 Two-phase melting problem with flux boundary condition
Consider a quasi one-dimensional computational domain $\Omega\in[0,0.1]\times[0,0.005]$ as illustrated in Fig. 15(A) [^8] This benchmark test is provided in IBAMR GitHub within the directory examples/phase_change/ex9.. The simulation employs a two-level mesh with coarse cell size of $\Delta x=\Delta y=0.005/32$. The simulation runs until $t=10$ s with a constant time step size of $\Delta t=10^{-3}$ s. Initially, the domain is fully occupied by the solid PCM with a temperature of $T_{i}=1500$ K. At the right boundary ($x=0.1$), a zero pressure (outflow) condition is used for the flow solver and an adiabatic (homogeneous Neumann) condition is used for the temperature $T$. At the left end of the domain ($x=0$), no-slip and inhomogeneous Neumann ${\left.-\kappa^{\rm L}\frac{{\rm d}T^{\rm L}_{\text{exact}}}{{\rm d}x}\right|_{x=0}}$ boundary conditions are employed for velocity and temperature, respectively. Here, $T^{\rm L}_{\text{exact}}$ denotes the analytical temperature of the liquid phase as derived in A. The top and bottom boundaries (in the $y$-direction) are periodic.
The schematic of the problem is depicted in Fig. 15(A). Since $R_{\rho}=\rho^{\rm S}/\rho^{\rm L}>1$, there is a volume expansion upon melting, which induces additional flow. As shown in A, the solid phase moves with a rigid body velocity $u^{\rm S}(t)$, while the liquid velocity $u^{\rm L}$ remains zero. To simulate this behavior, the solid fraction $\varphi_{\rm S}$ in the Carman-Kozeny drag coefficient is replaced with the liquid fraction $\varphi$. Consequently, the Carman-Kozeny drag force retards liquid phase motion. Gradient-based tagging is used to generate adaptively refined grids. Melting initiates due to the flux boundary condition and the liquid-solid interface moves rightwards over time until the PCM completely melts. The evolution of the liquid fraction and the adaptive mesh is shown in the left column of Fig. 15. We compare the numerical solutions (interface position $x*=s(t)$, velocity of the solid $u^{\rm S}(t)$, temperature distribution $T(x)$) with the analytical ones in Fig. 16. An excellent match is observed between the two. As shown in Fig. 15(B), a large spike is observed in the solid velocity at the initial time. This occurs because the interface velocity $u^{*}=\frac{{\rm d}s}{{\rm d}t}\rightarrow\infty$ (see Appendix A) at time $t=0^{+}$. Since the solid velocity is proportional to the interface speed, the CFD simulation produces a large solid velocity at the beginning. The new analytical and standard Stefan problem solutions are very similar for this problem. This is attributed to the 1D nature of the problem as well as the low density difference between the solid and liquid phases ($\approx 11\%$). Nevertheless, this amount of density variation may cause PCM free surfaces to rise or protrude in 2D or 3D problems [1].
@@
Figure 15: Stefan problem with melting: time evolution of the liquid fraction $\varphi$ and adaptive grids. For the left column plots, melting occurs due to the flux boundary condition at the left end of the domain. For the right column plots, flux boundary conditions are imposed through a volumetric term in the enthalpy equation. There is also a passive gas phase. Gradient-based tagging is used to generate AMR grids. Solid, liquid and gas regions are depicted in yellow, blue and white colors. The AMR grids are shown only at $t=5$ s.
@@ 7.3.2 Three-phase melting problem with a volumetric heat source
We now consider the three-phase version of the melting problem in which the heat flux boundary condition is imposed through a volumetric source term $Q_{\text{src}}$ [^9] This benchmark test is provided in IBAMR GitHub within the directory examples/phase_change/ex9.. The position of the heat source remains fixed at $x=0$, the initial location of the gas-PCM interface. Over time, the heat source strength varies (reduces). The PCM thermophysical properties remain the same as in the two-phase case.
@@
Due to the heat source, melting begins and progresses towards the right, as shown in Fig. 15. Comparison between numerical and analytical solutions is presented in Fig. 17. The numerical and analytical solutions are in excellent agreement. This simple test problem serves as a preliminary check to validate three phase codes aimed at modeling laser-induced PCM melting.
@@ 7.4 Porosity defect modeling
The final test case involves liquid, solid, and gas interactions. The goal is to simulate porosity defects during metal solidification, such as during metal 3D printing and casting. We consider a square computational domain $\Omega\in[0,L]^{2}$ of length $L=10^{-2}$ m.  [^10] This benchmark test is provided in IBAMR GitHub within the directory examples/phase_change/ex10. Fig. 18 depicts the initial phase distribution. The solid phase of the PCM occupies the region up to $y=0.1L$, the liquid phase of the PCM extends above it to $y=0.75L$, and a gas phase occupies the remainder. Three gas bubbles are placed within the liquid phase. The locations and radii of these bubbles can be found in Table 3.
@@
The simulation employs a two-level mesh with $N_{x0}\times N_{y0}=128^{2}$ coarse level cells, and a refinement ratio of 2 ($n_{\text{ref}}=2$). It runs until $t=1$ s with a uniform time step size of $3\times 10^{-6}$ s, which is determined by the explicit capillary time step size constraint. The PCM-gas interfacial region is refined using value-based ($|{\color[rgb]{0,0,0}\Phi}|_{i,j}\leq 2\Delta x_{0}$) tagging criterion. Gradient-based tagging refines the solid-liquid interfacial regions. As bubbles rise due to buoyancy, solidification occurs concurrently at the bottom wall due to the imposed temperature boundary condition. The movement of bubbles induces convective flow in the liquid below, which helps solidify it more quickly. After a while, the liquid-solid interface entraps the left and right bubbles, while the center bubble breaks off from the liquid-gas interface and enters the gas phase. Due to surface tension, the breaking of the liquid-gas interface initiates capillary waves, which eventually dissipate due to the combined effects of fluid viscosity and freezing. In the meantime, the two remaining bubbles are trapped within the solid. As a result of the slow solidification rate above the bubbles and the motion of the remaining untrapped portion, the left and right bubbles take on the shape of a pancake (flat on the top). Solidification continues until the liquid PCM fully solidifies. This is a shrinkage problem ($R_{\rho}=\rho^{\rm S}/\rho^{\rm L}>1$), so the top PCM-gas interface moves downwards due to volume changes, as shown in final Fig. 18(E). Upon completion of solidification, the interfaces remain stationary. At this point, the liquid phase has completely disappeared from the domain. The results of the previous Stefan problem section, as well as this section highlight the strength of our method to (i) handle the appearance and disappearance of phases in the simulation; (ii) properly account for volume changes in the material; and (iii) maintain equilibrium with no spurious phase generation in the absence of a phase change triggering mechanism.
@@
Figure 19: Percentage change in PCM mass $\mathcal{E_{M}}$ as a function of time for the metal solidification case with three gas bubbles. The coarsest grid $(N_{x}\times N_{y}=64^{2})$ uses a uniform time step size of $\Delta t=3\times 10^{-5}$ s. For each successively refined grid, the time step size is reduced by a factor of 3. The temperature interval considered in this case is $T^{\rm liq}-T^{\rm sol}=5$ K.
Figure 20: Percentage change in volume $\mathcal{E_{V}}$ as a function of time for the (A) left and (B) right bubbles for the metal solidification case with three gas bubbles. The coarsest grid $(N_{x}\times N_{y}=64^{2})$ uses a uniform time step size of $\Delta t=3\times 10^{-5}$ s. For each successively refined grid, the time step size is reduced by a factor of 3. The temperature interval considered in this case is $T^{\rm liq}-T^{\rm sol}=5$ K.
@@
Furthermore, we quantify the mass/volume conservation of our scheme in this case. The PCM mass is computed numerically as
@@
The relative change in the PCM mass is obtained using Eq. (43) and plotted against time for four uniform grids of size $N_{x}\times N_{y}=\{64^{2},128^{2},256^{2},512^{2}\}$ in Fig. 19. The coarsest grid $(N_{x}\times N_{y}=64^{2})$ uses a uniform time step size of $\Delta t=3\times 10^{-5}$ s. For each successively refined grid, the time step size is reduced by a factor of 3, respecting the explicit capillary time step size restriction.
It can be observed that there are two stages where PCM mass error increases over time. Initially, the PCM mass error increases rapidly until the center bubble is completely released into the air. Afterward, the PCM mass error remains constant for some time before growing again when the PCM-gas interface and liquid-solid interfaces approach each other. Around $t=0.3$ s, the gas-PCM and liquid-solid interfaces come close, resulting in an error in mass calculation because two different diffuse interfaces meet and equilibrate rapidly, causing the PCM to solidify completely. The non-linear convergence rate is attributed to errors associated with interface breaking and the interaction of two diffuse interfaces. The relative change in bubble volume is illustrated in Fig. 20. For all grids considered, the maximum volume change for the two bubbles is 4%. Given that the level set method leads to mass/volume loss, especially at coarse grid resolutions, this is an acceptable level of volume change.
We rerun the simulation after removing the center bubble from the domain to obtain meaningful convergence rates  [^11] Convergence rates are well-defined for smooth functions/dynamics.. This prevents interface breaking and ensures smooth dynamics. Fig. 21 shows the evolution of the interfaces. Solidification and bubble trapping dynamics are very similar to the three-bubble case discussed above. The same uniform grids are used to study mass errors as in the three bubble setup. The relative change in PCM mass is plotted against time in Fig. 22. It can be observed that as the grid cell size $\Delta$ decreases, the error in PCM mass, $\mathcal{E_{M}}$, also decreases until $t=0.3$ s. In the three bubble case, interface breaking causes PCM mass change errors to exhibit non-uniform convergence rates. In the two-bubble case without interface breaking, PCM mass change errors converge under grid refinement. Around $t=0.3$ s, the interface between gas-PCM and liquid-solid comes close to one another, leading to an increase in mass errors, which is similar to the three bubble setup. However, the % mass change of the PCM for all grids is quite low ($\approx 0.2\%$) compared to the bubble breaking case (roughly 1.4%). The relative change in bubble volume is illustrated in Fig. 23. For all grids considered, the maximum volume change for the two bubbles is 2% compared to 4% when bubble breaking is considered.